Sports shirts are frequently classified as S, M, L, XL for small, medium, large
and extra-large neck sizes. S fits a neck circumference of less than 37cm, M
fits between 37 and 40.5cm and L fits between 40.5cm and 44cm while XL fits
necks over 44cm in circumference. The neck circumference of adult males has
a normal distribution with  = 40cm and  = 2cm.
(i) What proportion of shirts should be manufactured in each category?
[6 marks]
(ii) If you wanted to define categories S, M, L, XL so that the categories contained
20%, 30%, 30% and 20% respectively, of the total population of
adult males, what neck sizes must you assign to each of the categories?
[6 marks]
1
Expert's answer
2021-09-06T15:03:15-0400

(i)

"S:"


"P(0<X<37)"

"=P(Z<\dfrac{37-40}{2})-P(Z\leq\dfrac{0-40}{2})"

"=P(Z<-1.5)-P(Z\leq-20)"

"\approx0.066807-0\approx0.0668"

"6.68\%"


"M:"


"P(37<X<40.5)"

"=P(Z<\dfrac{40.5-40}{2})-P(Z\leq\dfrac{37-40}{2})"

"=P(Z<0.25)-P(Z\leq-1.5)"

"\approx0.598706-0.066807\approx0.5319"

"53.19\%"



"L:"


"P(40.5<X<44)"

"=P(Z<\dfrac{44-40}{2})-P(Z\leq\dfrac{40.5-40}{2})"

"=P(Z<2)-P(Z\leq0.25)"

"\approx0.977250-0.598706\approx0.3785"

"37.85\%"



"XL:"


"P(X>44)"

"=1-P(Z\leq\dfrac{44-40}{2})"

"=1-P(Z\leq2)"

"\approx1-0.977250\approx0.0228"

"2.28\%"



"6.68+53.19+37.85+2.28=100"


(ii)

"S:"


"P(0<X<x_1)"

"=P(Z<\dfrac{x_1-40}{2})-P(Z\leq\dfrac{0-40}{2})"

"\approx P(Z<\dfrac{x_1-40}{2})-0\approx0.2"

"\dfrac{x_1-40}{2}\approx-0.8416""x_1\approx38.3"

"M:"


"P(38.3<X<x_2)"

"=P(Z<\dfrac{x_2-40}{2})-0.3=0.2"

"P(Z<\dfrac{x_2-40}{2})=0.5"

"\dfrac{x_2-40}{2}=0""x_2=40"

"L:"


"P(40<X<x_3)"

"=P(Z<\dfrac{x_3-40}{2})-0.5=0.3"

"P(Z<\dfrac{x_3-40}{2})=0.8"

"\dfrac{x_3-40}{2}=0.8416"




"x_3=41.7"

S fits a neck circumference of less than 38.3 cm, M fits between 38.3 cm and 40 cm and L fits between 40 cm and 41.7 cm while XL fits necks over 41.7 cm in circumference.